# Kinetic Interacting Particle Langevin Monte Carlo

### v2

@@ 2.3 MMLE via Interacting Particle Langevin Algorithm
These sampling algorithms, however, are not sufficient for parameter estimation. While sampling algorithms above provide an efficient way to sample from a given probability measure, they need to be supplemented by optimisation techniques when the target probability measure is indexed by a parameter $\theta$. As mentioned in the introduction, our main aim in this paper is to develop algorithms to solve the maximum marginal likelihood estimation problem MMLE which is given by
$\bar{\theta}_{\star}\in\operatornamewithlimits{argmax}_{\theta\in\mathbb{R}^{d_{\theta}}}\log p_{\theta}(y),$
for fixed $y\in\mathbb{R}^{d_{y}}$ and $p_{\theta}(y)=\int p_{\theta}(x,y){\mathrm{d}}x$. Therefore the MMLE problem involves the maximisation of an often intractable integral. Let $U(\theta,x)=-\log p_{\theta}(x,y)$ and note that the (intractable) $\theta$-gradient of $\log p_{\theta}(y)$ can be written as
$\nabla_{\theta}\log p_{\theta}(y)=\frac{-\int_{{\mathbb{R}}^{d_{x}}}\nabla_{\theta}U(\theta,x)p_{\theta}(x,y){\mathrm{d}}x}{p_{\theta}(y)}=-\int_{{\mathbb{R}}^{d_{x}}}\nabla_{\theta}U(\theta,x)p_{\theta}(x|y){\mathrm{d}}x,$ (6)
which is easy to prove (see, e.g., [29, Proposition D.4] and [3, Remark 1]). The expression in (6) is the motivation behind approximate schemes for EM, where one can first sample from $p_{\theta}(x|y)$ for fixed $\theta$ with an MCMC chain and then compute the gradient in (6) with these samples [26, 6]. However, these approaches require non-trivial assumptions on the step-size, as well as having a sample size that may need to grow to ensure that the gradient approximation does not incur asymptotic bias.
In contrast to this approach, [47] propose an IPS, the particle gradient descent (PGD) algorithm, to approximate the gradient of the $\theta$-dynamics by running independent particles to integrate out latent variables. Inspired by this, [3] attempt to solve the MMLE problem with a similar IPS, termed the interacting particle Langevin Algorithm (IPLA), which is a modification of PGD where $\theta$-dynamics contain a carefully scaled noise. This approach produces a system that is akin to the ULA and makes the theoretical analysis streamlined using the analysis produced for ULA, see, e.g., [31, 23, 32]. The proposed algorithm is a discretisation of a system of interacting Langevin SDEs which evolves in ${\mathbb{R}}^{d_{\theta}}\times{\mathbb{R}}^{Nd_{x}}$ where $N$ distinct particles are retained for latent variables. More precisely, IPLA recursions are based on the SDE:
@@
$\mathrm{d}{\mathbf{X}}_{t}^{i}$ $=-\nabla_{x}U(\bm{\theta}_{t},{\mathbf{X}}^{i}_{t})\mathrm{d}t+\sqrt{2}\mathrm{d}{\mathbf{B}}_{t}^{i},\qquad\text{for }i\in[N].$ (8)
where $({\mathbf{B}}_{t}^{0})_{t\geq 0}$ is a Brownian motion evolving on ${\mathbb{R}}^{d_{\theta}}$ and $({\mathbf{B}}_{t}^{i})_{t\geq 0}$ for $i\in[N]$ are Brownian motions evolving on $\mathbb{R}^{d_{x}}$. Intuitively, it can be seen that the drift term for the $\theta$-dynamics can be interpreted as the empirical approximation of the “true” drift (6) with $N$ particles drawn from a measure drifting towards $p_{\theta}(x|y)$. In this case, [3] observe that the stationary measure of this system is given as
$\pi^{N}({\mathrm{d}}\theta,{\mathrm{d}}x_{1},\dots,{\mathrm{d}}x_{N})\propto\exp\left(-\sum_{i=1}^{N}U(\theta,x_{i})\right){\mathrm{d}}\theta{\mathrm{d}}x_{1}\dots\mathrm{d}x_{N}.$
It is easy to show that, the $\theta$-marginal of this measure concentrates on the MMLE $\bar{\theta}_{\star}$ as $N$ grows, similar to annealing techniques developed in traditional optimisation [41]. The authors of [3] identify the convergence rate to the joint stationary measure, as well as, an error bound for the discretisation error, from which an error can be determined between the $\theta$-iterate of the algorithm and the MMLE. With this nonasymptotic analysis, [3] identify parameters needed to implement the algorithm and introduce an order of convergence guarantee in $W_{2}$. The algorithm is empirically shown to be competitive with similar methods, such as SOUL [26] and the PGD proposed in [47], while also amenable to a streamlined theoretical analysis akin to ULA. Furthermore, the similarity between IPLA and ULA lays the foundation for several avenues of further exploration, such as considering a kinetic Langevin algorithm and exploring other numerical discretizations of the proposed SDEs, as we aim to do in this paper.
@@ 3 Kinetic Interacting Particle Langevin Monte Carlo
Our goal is to combine the advantages of underdamped Langevin diffusions with the interacting particle systems proposed by [3, 47, 49], to estimate the MMLE $\bar{\theta}_{\star}$. In this section we present a diffusion and two algorithms to this end. Indeed, this diffusion will be an underdamped version of the system of overdamped Langevin diffusions of the IPLA SDEs given in (7)–(8).
3.1 Kinetic Interacting Particle Langevin Diffusion
We introduce the KIPLD system, the underdamped counterpart of the IPLA diffusion given. More precisely, the KIPLD diffusion evolves on ${\mathbb{R}}^{2d_{\theta}}\times{\mathbb{R}}^{2Nd_{x}}$ and given by
@@ 3.1 Kinetic Interacting Particle Langevin Diffusion
$\mathrm{d}{\mathbf{V}}^{x_{i}}_{t}$ $=-\gamma{\mathbf{V}}_{t}^{x_{i}}\mathrm{d}t-\nabla_{x}U(\bm{\theta}_{t},{\mathbf{X}}_{t}^{i})\mathrm{d}t+\sqrt{2\gamma}\mathrm{d}{\mathbf{B}}_{t}^{i},\qquad\text{for }i\in[N]$
where $({\mathbf{B}}_{t}^{i})_{t\geq 0}$ for $i\in[N]$ is a family of ${\mathbb{R}}^{d_{x}}$-valued Brownian motions and $({\mathbf{B}}_{t}^{0})_{t\geq 0}$ is an $\mathbb{R}^{d_{\theta}}$-valued Brownian motion. For notational convenience we will say that the particles and momenta have combined dimension of $d_{z}=d_{\theta}+Nd_{x}$ each. Similarly to the algorithm proposed by [3], we have a system of $N$ particles for the empirical approximation of latent variable distribution. This system corresponds to relaxing the limit on $\gamma$, the “friction” parameter, for the algorithm proposed by [3]. Adding the friction hyper-parameter $\gamma$ should allow for improved regularisation and dampening momentum effects, as discussed in [51]. We will note some of this behaviour in the experimental section, where the momentum effect can be observed and tuned to produce optimal convergence behaviour.
@@
3.2 Kinetic Interacting Particle Langevin Monte Carlo Methods
We now introduce two numerical integrators for the KIPLD: firstly, we consider an Exponential Integrator, as discussed in [25]; secondly, a splitting scheme is applied, as described in [53]. These algorithms for kinetic Langevin samplers have two parameters, namely, the step-size $\eta>0$, which assumed to be small, and the friction coefficient $\gamma>0$, determining momentum effects.
3.2.1 Exponential Integrator (KIPLMC1)
In this section, we consider a discretisation similar to the one employed by [20], which is an exponential integrator tailored to the kinetic Langevin diffusion. This discretisation is a variant of the Hamiltonian Monte Carlo which has improved convergence properties when compared to the Langevin MCMC proposed by [32]. Specifically, while the standard overdamped Langevin methods (e.g. ULA) requires $\widetilde{\mathcal{O}}(d/\varepsilon^{2})$ number of steps [^3] The notation $\widetilde{\mathcal{O}}$ supresses the logarithmic dependence to $d$ and $\varepsilon$. to attain $\varepsilon$ error in Wasserstein-2 distance, it is shown in [20, Theorem 1] that the underdamped Langevin MCMC using the exponential integrator requires $\widetilde{\mathcal{O}}(\sqrt{d}/\varepsilon)$ steps to attain the same error rate. This method is proposed for sampling from underdamped Langevin diffusions, as, unlike Euler-Maruyama, this approach is able to exploit the higher order of convergence of the underlying process [25]. The bounds on the convergence rate for this algorithm are further improved in [25] with a tightened the bound with respect to the condition number $L/\mu$ and improve sensitivity to the initial choice.
Motivated by this we aim at applying this scheme to the diffusion KIPLD and seek to recover improved results (see Section 4.4 for details). Following [25], we begin by defining functions
$\psi_{0}(t)$ $=e^{-\gamma t}$
$\psi_{1}(t)$ $=\int_{0}^{t}e^{-\gamma s}{\mathrm{d}}s=\frac{1}{\gamma}(e^{-\gamma t}-1)$
$\psi_{2}(t)$ $=\frac{1}{\gamma}\int_{0}^{t}(e^{-\gamma s}-1){\mathrm{d}}s=\frac{1}{\gamma^{2}}(e^{-\gamma t}-1)-\frac{t}{\gamma}.$
These terms emerge from the Itô formula used to estimate the expectations and their covariances (see [20, Lemma 11] for a full derivation). The pairs $\varepsilon_{n}^{i}$, $\varepsilon_{n}^{i,\prime}$, for $i\in[N]$, are i.i.d. standard normal Gaussians with pair-wise covariance matrix
$C=\int_{0}^{\eta}\begin{pmatrix}\psi_{0}(t)^{2}&\psi_{0}(t)\psi_{1}(t)\\ \psi_{0}(t)\psi_{1}(t)&\psi_{1}(t)^{2}\end{pmatrix}\mathrm{d}t.$ (9)
This covariance matrix is the covariance between the dynamics of the position particle and the momentum particle, as shown in [20, Lemma 11]. The scheme is recovered by updating the particle positions with the expectations of the particles conditioned on the previous time-step and adding Gaussian noise with covariance $C$. Thus, the scheme produces the same first and second momenta as the diffusion we seek to target. Formally, we define the KIPLMC1 scheme as
$\theta_{n+1}$ $=\theta_{n}+\psi_{1}(\eta)V^{\theta}_{n}-\frac{\psi_{2}(\eta)}{N}\sum_{i=1}^{N}\nabla_{\theta}U(\theta_{n},X^{i}_{n})+\sqrt{\frac{2\gamma}{N}}\varepsilon^{0,\prime}_{n}$ (KIPLMC1)
$X^{i}_{n+1}$ $=X^{i}_{n}+\psi_{1}(\eta)V^{x_{i}}_{n}-\psi_{2}(\eta)\nabla_{x}U(\theta_{n},X^{i}_{n})+\sqrt{2\gamma}\varepsilon_{n}^{i,\prime}$
$V^{\theta}_{n+1}$ $=\psi_{0}(\eta)V^{\theta}_{n}-\frac{\psi_{1}(\eta)}{N}\sum_{i=1}^{N}\nabla_{\theta}U(\theta_{n},X^{i}_{n})+\sqrt{\frac{2\gamma}{N}}\varepsilon^{0}_{n}$
$V^{x_{i}}_{n+1}$ $=\psi_{0}(\eta)V^{x_{i}}_{n}-\psi_{1}(\eta)\nabla_{x}U(\theta_{n},X^{i}_{n})+\sqrt{2\gamma}\varepsilon^{i}_{n}$
for $i\in[N]$, where $\varepsilon_{n}$ and $\varepsilon^{\prime}_{n}$ are standard Gaussians with pairwise covariance given by $C$. The scheme converges for certain choices of $\eta$, depending on the assumptions set on $U$. The full algorithm is given in Algorithm 1.
Algorithm 1 KIPLMC1 Algorithm
1:$\gamma,\eta>0$, $N,K\in{\mathbb{N}}$

### v3

@@ 2.2 MMLE via Interacting Particle Langevin Algorithm
Consider the MMLE problem given in (1), that is, to find the parameter $\theta$ that maximises the marginal likelihood $p_{\theta}(y)$. As summarised in the introduction, this problem is often solved with the EM algorithm, typically implemented with a double-loop structure consisting of MCMC techniques for the E-step and numerical optimisation techniques for the M-step [6, 21].
In contrast to this approach, Kuntz et al. [35] propose an interacting particle system, the particle gradient descent (PGD) algorithm, to approximate the gradient of the $\theta$-dynamics by running independent particles to integrate out latent variables. Inspired by this, Akyildiz et al. [1] attempt to solve the MMLE problem with a similar system, termed the interacting particle Langevin algorithm (IPLA), which is a modification of PGD where $\theta$-dynamics contain a carefully scaled noise. This approach produces a system that is akin to the ULA and makes the theoretical analysis streamlined using the analysis produced for ULA [20, 25]. The proposed algorithm is a discretisation of a system of interacting Langevin SDEs which evolves in ${\mathbb{R}}^{d_{\theta}}\times{\mathbb{R}}^{Nd_{x}}$ where $N$ distinct particles are retained for latent variables. More precisely, IPLA recursions are based on the SDE:
@@
$\mathrm{d}{\mathbf{X}}_{t}^{i}$ $=-\nabla_{x}U(\bm{\theta}_{t},{\mathbf{X}}^{i}_{t})\mathrm{d}t+\sqrt{2}\mathrm{d}{\mathbf{B}}_{t}^{i},$ (6)
for $i\in[N]$, where $({\mathbf{B}}_{t}^{0})_{t\geq 0}$ is a Brownian motion evolving on ${\mathbb{R}}^{d_{\theta}}$ and $({\mathbf{B}}_{t}^{i})_{t\geq 0}$ for $i\in[N]$ are Brownian motions evolving on $\mathbb{R}^{d_{x}}$. In this case, Akyildiz et al. [1] observe that the $\theta$-marginal of the stationary measure of this system takes the form $\pi_{\Theta}({\mathrm{d}}\theta)\propto\exp(N\log p_{\theta}(y)){\mathrm{d}}\theta$ which concentrates on the MMLE $\bar{\theta}_{\star}$ as $N$ grows, where $N$ plays the role of inverse temperature [31]. The authors then identify the convergence rate to the joint stationary measure, as well as an error bound for the discretisation error, from which an error can be determined between the $\theta$-iterate of the algorithm and the MMLE.
@@ 3 Kinetic Interacting Particle Langevin Monte Carlo
To develop our methodology, we start with the following diffusion process:
@@
$\mathrm{d}{\mathbf{V}}^{x_{i}}_{t}$ $=-\gamma{\mathbf{V}}_{t}^{x_{i}}\mathrm{d}t-\nabla_{x}U(\bm{\theta}_{t},{\mathbf{X}}_{t}^{i})\mathrm{d}t+\sqrt{2\gamma}\mathrm{d}{\mathbf{B}}_{t}^{i},$
for $i\in[N]$, where $\{({\mathbf{B}}_{t}^{i})_{t\geq 0}\}_{i\in[N]}$ is a family of ${\mathbb{R}}^{d_{x}}$-valued Brownian motions and $({\mathbf{B}}_{t}^{0})_{t\geq 0}$ is an $\mathbb{R}^{d_{\theta}}$-valued Brownian motion. This diffusion has the property that its $\theta$-marginal at stationarity concentrates around the global minimisers of $\log p_{\theta}(y)$ (see Section 4).
Next, we introduce two numerical integrators for the KIPLD: firstly, we consider an Exponential Integrator, as discussed in Dalalyan and Riou-Durand [19]; secondly, a splitting scheme is applied, as described in Monmarché [41].
3.1 Exponential Integrator (KIPLMC1)
In order to derive the exponential integrator discretisation of KIPLD, following Dalalyan and Riou-Durand [19], we begin by defining the functions
$\psi_{0}^{t}=e^{-\gamma t},\quad\psi_{1}^{t}=\int_{0}^{t}e^{-\gamma s}{\mathrm{d}}s=\frac{1}{\gamma}(1-e^{-\gamma t}),\quad\psi_{2}^{t}=\int_{0}^{t}\psi_{1}^{s}{\mathrm{d}}s=\frac{1}{\gamma^{2}}(e^{-\gamma t}-1)+\frac{t}{\gamma}.$
The exponential integrator is obtained by freezing the nonlinear drift terms $\nabla_{\theta}U(\theta_{t},X_{t}^{i})$ and $\nabla_{x}U(\theta_{t},X_{t}^{i})$ at the beginning of each time step and solving exactly the resulting linear stochastic differential equation.
More precisely, let $t_{n}=n\eta$. Over the interval $t_{n},t_{n+1}]$ we approximate the dynamics (KIPLD) by freezing the gradients at $(\theta_{n},X_{n}^{i})$,
$F_{n}^{\theta}=\frac{1}{N}\sum_{i=1}^{N}\nabla_{\theta}U(\theta_{n},X_{n}^{i}),\qquad F_{n}^{x_{i}}=\nabla_{x}U(\theta_{n},X_{n}^{i}),$
which yields the linear system
$\mathrm{d}\bm{\theta}_{t}$ $={\mathbf{V}}_{t}^{\theta}\mathrm{d}t,$
$\mathrm{d}{\mathbf{X}}_{t}^{i}$ $={\mathbf{V}}_{t}^{x_{i}}\mathrm{d}t,$
$\mathrm{d}{\mathbf{V}}_{t}^{\theta}$ $=-\gamma{\mathbf{V}}_{t}^{\theta}\mathrm{d}t-F_{n}^{\theta}\mathrm{d}t+\sqrt{\frac{2\gamma}{N}}\mathrm{d}{\mathbf{B}}_{t}^{0},$
$\mathrm{d}{\mathbf{V}}_{t}^{x_{i}}$ $=-\gamma{\mathbf{V}}_{t}^{x_{i}}\mathrm{d}t-F_{n}^{x_{i}}\mathrm{d}t+\sqrt{2\gamma}\mathrm{d}{\mathbf{B}}_{t}^{i},$
for $i\in[N]$, with initial condition
$(\bm{\theta}_{t_{n}},{\mathbf{X}}_{t_{n}}^{i},{\mathbf{V}}_{t_{n}}^{\theta},{\mathbf{V}}_{t_{n}}^{x_{i}})=(\theta_{n},X_{n}^{i},V_{n}^{\theta},V_{n}^{x_{i}}).$
Since the system is linear with constant coefficients, it can be solved explicitly. Using an integrating factor for the velocity equation gives, for $s\in[0,\eta]$,
$V_{t_{n}+s}^{\theta}$ $=\psi_{0}^{s}V_{n}^{\theta}-\psi_{1}^{s}F_{n}^{\theta}+\sqrt{\frac{2\gamma}{N}}\int_{0}^{s}\psi_{0}^{\,s-\tau}\,\mathrm{d}{\mathbf{B}}_{t_{n}+\tau}^{0},$
$V_{t_{n}+s}^{x_{i}}$ $=\psi_{0}^{s}V_{n}^{x_{i}}-\psi_{1}^{s}F_{n}^{x_{i}}+\sqrt{2\gamma}\int_{0}^{s}\psi_{0}^{\,s-\tau}\,\mathrm{d}{\mathbf{B}}_{t_{n}+\tau}^{i}.$
Algorithm 1 KIPLMC1
@@ 3.1 Exponential Integrator (KIPLMC1)
2:Compute the covariance matrix $C$ given in (7) and its Cholesky decomposition $C=LL^{\top}$.
3:Draw $\theta_{0},X^{i}_{0},V^{\theta}_{0},V^{x_{i}}_{0}$, for $i\in[N]$
4:for n=0:M1n=0:M-1 do
5:  Draw: $\xi_{n}^{i},\xi_{n}^{i,\prime}$ from a standard Gaussian (in $\mathbb{R}^{d_{\theta}}$ for $i=0$ and $\mathbb{R}^{d_{x}}$ otherwise)
$\varepsilon_{n}^{i}=L_{11}\xi_{n}^{i},\qquad\varepsilon_{n}^{i,\prime}=L_{21}\xi_{n}^{i}+L_{22}\xi_{n}^{i,\prime}$
for $i=0,\ldots,N$ where $\xi_{n}^{0}$ and $\xi_{n}^{0,\prime}$ are standard $d_{x}$ and $d_{\theta}$ dimensional Normal random variables.
6:  Compute the average gradient of the potential:
$F_{n}^{\theta}=\frac{1}{N}\sum_{i=1}^{N}\nabla_{\theta}U(\theta_{n},X^{i}_{n}),\qquad F_{n}^{x_{i}}=\nabla_{x}U(\theta_{n},X^{i}_{n}).$
Update the parameters, particles and their momenta:
$\theta_{n+1}$ $=\theta_{n}+\psi_{1}^{\eta}V^{\theta}_{n}-{\psi_{2}^{\eta}}F_{n}^{\theta}+\sqrt{\frac{2\gamma}{N}}\varepsilon^{0,\prime}_{n}$
$X^{i}_{n+1}$ $=X^{i}_{n}+\psi_{1}^{\eta}V^{x_{i}}_{n}-\psi_{2}^{\eta}F_{n}^{x_{i}}+\sqrt{2\gamma}\varepsilon_{n}^{i,\prime}$ (KIPLMC1)
$V^{\theta}_{n+1}$ $=\psi_{0}^{\eta}V^{\theta}_{n}-\psi_{1}^{\eta}F_{n}^{\theta}+\sqrt{\frac{2\gamma}{N}}\varepsilon^{0}_{n}$
$V^{x_{i}}_{n+1}$ $=\psi_{0}^{\eta}V^{x_{i}}_{n}-\psi_{1}^{\eta}F_{n}^{x_{i}}+\sqrt{2\gamma}\varepsilon^{i}_{n}$
for $i\in[N]$.
7:Output: $(\theta_{M},X^{1}_{M},\ldots,X^{N}_{M})$
Integrating once more yields the updates for the positions
$\theta_{t_{n}+s}$ $=\theta_{n}+\psi_{1}^{s}V_{n}^{\theta}-\psi_{2}^{s}F_{n}^{\theta}+\sqrt{\frac{2\gamma}{N}}\int_{0}^{s}\psi_{1}^{\,s-\tau}\,\mathrm{d}{\mathbf{B}}_{t_{n}+\tau}^{0},$
$X_{t_{n}+s}^{i}$ $=X_{n}^{i}+\psi_{1}^{s}V_{n}^{x_{i}}-\psi_{2}^{s}F_{n}^{x_{i}}+\sqrt{2\gamma}\int_{0}^{s}\psi_{1}^{\,s-\tau}\,\mathrm{d}{\mathbf{B}}_{t_{n}+\tau}^{i}.$
Evaluating these expressions at $s=\eta$ gives the one-step update of the discretised process. The stochastic integrals appearing above are jointly Gaussian random variables. For each $i\in\{0,\ldots,N\}$ we define
$\varepsilon_{n}^{i}=\int_{0}^{\eta}\psi_{0}^{\,\eta-s}\,\mathrm{d}B_{t_{n}+s}^{i},\quad\varepsilon_{n}^{i,\prime}=\int_{0}^{\eta}\psi_{1}^{\,\eta-s}\,\mathrm{d}B_{t_{n}+s}^{i}.$
Now note that $(\varepsilon_{n}^{0},\varepsilon_{n}^{0,\prime})\in\mathbb{R}^{d_{\theta}}\times\mathbb{R}^{d_{\theta}}$ and $(\varepsilon_{n}^{i},\varepsilon_{n}^{i,\prime})\in\mathbb{R}^{d_{x}}\times\mathbb{R}^{d_{x}}$ for $i=1,\ldots,N$. Consider each coordinate of these pairs:
$\varepsilon_{n,j}^{i}=\int_{0}^{\eta}\psi_{0}^{\,\eta-s}\,\mathrm{d}B_{t_{n}+s,j}^{i},\quad\varepsilon_{n,j}^{i,\prime}=\int_{0}^{\eta}\psi_{1}^{\,\eta-s}\,\mathrm{d}B_{t_{n}+s,j}^{i}.$
For every $(i,j)$, using Itô isometry [42], we can obtain $C=\mathrm{Cov}(\varepsilon_{n,j}^{i},\varepsilon_{n,j}^{i,\prime})$ as
$C=\int_{0}^{\eta}\begin{pmatrix}(\psi_{0}^{t})^{2}&\psi_{0}^{t}\psi_{1}^{t}\\ \psi_{0}^{t}\psi_{1}^{t}&(\psi_{1}^{t})^{2}\end{pmatrix}\mathrm{d}t.$ (7)
We note that $C\in\mathbb{R}^{2\times 2}$ and denote its Cholesky decomposition by $C=LL^{\top}$. Hence the Gaussian pair $(\varepsilon_{n,j}^{i},\varepsilon_{n,j}^{i,\prime})$ can be sampled by setting
$\begin{pmatrix}\varepsilon_{n,j}^{i}\\ \varepsilon_{n,j}^{i,\prime}\end{pmatrix}$ $=L\begin{pmatrix}\xi_{n,j}^{i}\\ \xi_{n,j}^{i,\prime}\end{pmatrix}$,
where $\xi_{n,j}^{i}$ and $\xi_{n,j}^{i,\prime}$ are independent standard Gaussian random variables.
Substituting these expressions for $s=\eta$ yields the KIPLMC1 scheme. The full algorithm is given in Algorithm 1 and a more explicit derivation in Appendix B.
3.2 A Splitting Scheme (KIPLMC2)
We next introduce KIPLMC2, an adaptation of the underdamped Langevin sampler introduced by Horowitz [30] and extensively studied by Monmarché [41]. This splitting scheme is termed OBABO and it is a second order scheme, only requiring the first order derivatives of $U$. The idea is to split the numerical scheme into simpler subflows that can be solved explicitly. More precisely, in our KIPLD case, we decompose the flow into a Ornstein-Uhlenbeck step (O), a velocity drift step (B) and a particle update step (A).
Algorithm 2 KIPLMC2 Algorithm
[…]